OpenStudy (anonymous):

Simplify (3ab^7)^2. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

square each term

jimthompson5910 (jim_thompson5910):

and use the idea that (a^b)^c = a^(b*c)

OpenStudy (anonymous):

We have : \[(3ab^7)^2=3^2a^2\left(b^7\right)^2=3^2a^2b^{7\times2}=3^2a^2b^{14}\]

OpenStudy (anonymous):

I have two options that look like yours, @Noura11 but instead of a 3 at the front, it's a 6 or a 9.

jimthompson5910 (jim_thompson5910):

3^2 is out front, not just 3

OpenStudy (anonymous):

3^2=3*3=9

OpenStudy (anonymous):

Oops >.< Thanks guys.